OpenStudy (anonymous):

The Shelby University rose garden is a rectangular shape with an area of 192 square feet. The dimensions of the garden are x^2 feet wide by x^2-61 feet long. How wide is the garden ? A.3 feet B.8 feet c.64 feet D.131 feet

OpenStudy (anonymous):

Area=Length*breath \[192=x^2(x^2-61)\] Put t=x^2 and solve the quadratic

OpenStudy (anonymous):

i think i got it . Thank You ! (: